# InternVL3.5: Advancing Open-Source Multimodal Models in Versatility, Reasoning, and Efficiency

### v1

@@ 1 Introduction
The recent trend of Multimodal Large Language Models (MLLMs) qwen2.5 ; team2025keye_vl ; hong2025glm_v_thinking has gone beyond simple multimodal understanding and gradually focused on more general, complex, and realistic tasks such as text-related tasks aime2024 ; chen2021evaluating ; hendrycks2020measuring ; Zhang2023gaokao ; huang2023ceval ; rein2024gpqa , reasoning tasks yang2024think ; wang2024mathvision ; lu2023mathvista ; yue2023mmmu ; he2024olympiadbench and agentic tasks qin2025ui ; hu2025os_agents ; xie2024osworld ; zhang2025guimid ; yang2025zerogui ; sun2025scienceboard ; jia2025omnispatial ; yang2024think ; gong2025space10 . In these aspects, commercial models have created huge gaps with current open-source models, as shown in Table 2. Thereby, recent open-source efforts meng2025mm_eureka ; hong2025glm_v_thinking ; team2025keye_vl ; wang2025skywork_r1v2 aim to explore advanced reinforcement learning (RL) methods to mitigate the gap and pursue higher multimodal intelligence. However, despite much effort in RL algorithms shao2024deepseekmath ; zheng2025gspo ; meng2025mm_eureka ; cui2025prime ; cui2025entropy_mechanism ; yue2025does ; vl-rethinker ; xi2025bmmr and verifiers wang2025visualprm ; wang2023mathshepherd ; luo2024omegaprm , a stable, effective, and scalable reinforcement learning framework for MLLMs still remains an open problem in the community. Furthermore, the growth of multimodal capabilities, e.g., long visual context and high-resolution understanding zhu2025internvl3 ; llava_uhd_v2 ; luo2024llava_hr ; wang2025parameter ; bai2025qwen2_5 , often comes with ever increasing computational costs, which have become a crucial bottleneck of real-world applications.
@@ 3.1 Overall Comparison with Other Advanced MLLMs
Table 2 presents a comprehensive evaluation of InternVL3.5’s performance across 36 benchmarks categorized into four key multimodal task types: (1) General Tasks: MMStar chen2024mmstar , MMVet yu2023mmvet , MMBench V1.1 (en) liu2023mmbench , MTVQA tang2024mtvqa , AI2D kembhavi2016ai2d , OCRBench liu2023ocrbench , WildVision lu2024wildvision , MME-RealWorld (en) zhang2024mme , HallusionBench guan2023hallusionbench , MVBench li2024mvbench , VideoMME fu2024video , MLVU MLVU , LVBench wu2024longvideobench ; (2) Reasoning Tasks: MMMU yue2023mmmu , MathVista lu2023mathvista , MathVision lu2023mathvista , MathVerse zhang2024mathverse , DynaMath zou2024dynamath , WeMath qiao2024wemath , OlympiadBench he2024olympiadbench , LogicVista xiao2024logicvista ; (3) Text-Centric Tasks: MATH500 lightman2023prm800k , AIME24 aime2024 , AIME25 aime2025 , GPQA rein2024gpqa , MMLU-Pro li2023cmmlu , GAOKAO Zhang2023gaokao , IFEval zhou2023instruction ; (4) Agentic Tasks: SGP-Bench qiu2024can , ScreenSpot cheng2024seeclick , ScreenSpot-v2 wu2024atlas , OSWorld-G xie2024osworld , VSI-Bench yang2024think , ERQA erqa , SpaCE-10 gong2025space10 .
@@
These results highlight InternVL3.5’s strong capabilities across diverse tasks. For general multimodal tasks, InternVL3.5 demonstrates leading performance among open-source models on general multimodal understanding (e.g., MMVet), multilingual (e.g., MTVQA), OCR (e.g., OCRBench), real-world (e.g., MME-RealWorld), and video (e.g., LVBench) benchmarks. It even achieves a similar overall score as GPT-5 gpt5 (74.1 vs. 74.0), the state-of-the-art closed-source MLLM. Nevertheless, some tasks like HallusionBench still pose challenges, indicating the need for further refinements.
@@ 3.3 OCR, Chart, and Document Understanding
As shown in Table 4, InternVL3.5 achieves competitive results on these benchmarks, outperforming other open-source and closed-source models. At the lightweight scale, InternVL3.5 demonstrates significant potential. For instance, InternVL3.5-2B attains an overall average score of 76.5 across nine benchmarks, surpassing InternVL3-2B of similar size, which scores 74.7. Specifically, on DocVQA, InfoVQA, and SEED-2-Plus, InternVL3.5-2B achieves scores of 89.4, 68.2, and 68.0, respectively, compared to 88.3, 66.1, and 64.6 for InternVL3-2B, highlighting InternVL3.5’s strong vision-language understanding capabilities.
Moreover, as the model scale increases, InternVL3.5 continues to deliver enhanced performance in vision-language understanding. Across the same nine benchmarks, InternVL3.5-4B, InternVL3.5-20B-A4B, InternVL3.5-14B, InternVL3.5-30B-A3B, and InternVL3.5-38B achieve overall average scores of 79.7, 81.7, 82.0, 83.6, and 84.4, respectively, demonstrating consistent improvements with larger model sizes. On AI2D, InternVL3.5-2B, InternVL3.5-4B, InternVL3.5-14B and InternVL3.5-38B obtain scores of 78.8/89.1, 82.6/92.3, 85.1/93.3, and 87.8/95.1, respectively, indicating a consistent upward trend across larger models. On ChartQA, InternVL3.5-4B shows a notable improvement in chart understanding, increasing from 80.7 (InternVL3.5-2B) to 86.0, reflecting a significant gain in visual reasoning capability. Similarly, on CharXiv and VCR-EN-Easy, larger model variants also achieve measurable improvements in text and document understanding.

### v2

@@ 1 Introduction
The recent trend of Multimodal Large Language Models (MLLMs) qwen2.5 ; team2025keye_vl ; hong2025glm_v_thinking has gone beyond simple multimodal understanding and gradually focused on more general, complex, and realistic tasks such as text-related tasks aime2024 ; chen2021evaluating ; hendrycks2020measuring ; Zhang2023gaokao ; huang2023ceval ; rein2024gpqa , reasoning tasks yang2024think ; wang2024mathvision ; lu2023mathvista ; yue2023mmmu ; he2024olympiadbench and agentic tasks qin2025ui ; hu2025os_agents ; xie2024osworld ; zhang2025guimid ; yang2025zerogui ; sun2025scienceboard ; jia2025omnispatial ; yang2024think ; gong2025space10 . In these aspects, commercial models have created huge gaps with current open-source models, as shown in Table 2. Thereby, recent open-source efforts meng2025mm_eureka ; hong2025glm_v_thinking ; team2025keye_vl ; wang2025skywork_r1v2 aim to explore advanced reinforcement learning (RL) methods to mitigate the gap and pursue higher multimodal intelligence. However, despite much effort in RL algorithms shao2024deepseekmath ; zheng2025gspo ; meng2025mm_eureka ; cui2025prime ; cui2025entropy_mechanism ; yue2025does ; vl-rethinker ; xi2025bmmr and verifiers wang2025visualprm ; wang2023mathshepherd ; luo2024omegaprm , a stable, effective, and scalable reinforcement learning framework for MLLMs still remains an open problem in the community. Furthermore, the growth of multimodal capabilities, e.g., long visual context and high-resolution understanding zhu2025internvl3 ; llava_uhd_v2 ; luo2024llava_hr ; wang2025parameter ; bai2025qwen2_5 ; wang2025lvbenchextremelongvideo , often comes with ever increasing computational costs, which have become a crucial bottleneck of real-world applications.
@@ 3.1 Overall Comparison with Other Advanced MLLMs
Table 2 presents a comprehensive evaluation of InternVL3.5’s performance across 35 benchmarks categorized into four key multimodal task types: (1) General Tasks: MMStar chen2024mmstar , MMVet yu2023mmvet , MMBench V1.1 (en) liu2023mmbench , MTVQA tang2024mtvqa , AI2D kembhavi2016ai2d , OCRBench liu2023ocrbench , WildVision lu2024wildvision , MME-RealWorld (en) zhang2024mme , HallusionBench guan2023hallusionbench , MVBench li2024mvbench , VideoMME fu2024video , MLVU MLVU , (2) Reasoning Tasks: MMMU yue2023mmmu , MathVista lu2023mathvista , MathVision lu2023mathvista , MathVerse zhang2024mathverse , DynaMath zou2024dynamath , WeMath qiao2024wemath , OlympiadBench he2024olympiadbench , LogicVista xiao2024logicvista ; (3) Text-Centric Tasks: MATH500 lightman2023prm800k , AIME24 aime2024 , AIME25 aime2025 , GPQA rein2024gpqa , MMLU-Pro li2023cmmlu , GAOKAO Zhang2023gaokao , IFEval zhou2023instruction ; (4) Agentic Tasks: SGP-Bench qiu2024can , ScreenSpot cheng2024seeclick , ScreenSpot-v2 wu2024atlas , OSWorld-G xie2024osworld , VSI-Bench yang2024think , ERQA erqa , SpaCE-10 gong2025space10 .
@@
These results highlight InternVL3.5’s strong capabilities across diverse tasks. For general multimodal tasks, InternVL3.5 demonstrates leading performance among open-source models on general multimodal understanding (e.g., MMVet), multilingual (e.g., MTVQA), OCR (e.g., OCRBench), real-world (e.g., MME-RealWorld), and video (e.g., LongVideoBench) benchmarks. It even achieves a similar overall score as GPT-5 gpt5 (74.1 vs. 74.0), the state-of-the-art closed-source MLLM. Nevertheless, some tasks like HallusionBench still pose challenges, indicating the need for further refinements.
@@ 3.3 OCR, Chart, and Document Understanding
As shown in Table 4, InternVL3.5 achieves competitive results on these benchmarks, outperforming other open-source and closed-source models. At the lightweight scale, InternVL3.5 demonstrates significant potential. For instance, InternVL3.5-2B attains an overall average score of 76.7 across nine benchmarks, surpassing InternVL3-2B of similar size, which scores 74.7. Specifically, on DocVQA, InfoVQA, and SEED-2-Plus, InternVL3.5-2B achieves scores of 89.4, 70.8, and 68.0, respectively, compared to 88.3, 66.1, and 64.6 for InternVL3-2B, highlighting InternVL3.5’s strong vision-language understanding capabilities.
Moreover, as the model scale increases, InternVL3.5 continues to deliver enhanced performance in vision-language understanding. Across the same nine benchmarks, InternVL3.5-4B, InternVL3.5-20B-A4B, InternVL3.5-14B, InternVL3.5-30B-A3B, and InternVL3.5-38B achieve overall average scores of 80.0, 81.7, 82.0, 83.9, and 84.6, respectively, demonstrating consistent improvements with larger model sizes. On AI2D, InternVL3.5-2B, InternVL3.5-4B, InternVL3.5-14B and InternVL3.5-38B obtain scores of 78.8/89.1, 82.6/92.3, 85.1/93.3, and 87.8/95.1, respectively, indicating a consistent upward trend across larger models. On ChartQA, InternVL3.5-4B shows a notable improvement in chart understanding, increasing from 80.7 (InternVL3.5-2B) to 86.0, reflecting a significant gain in visual reasoning capability. Similarly, on CharXiv and VCR-EN-Easy, larger model variants also achieve measurable improvements in text and document understanding.